The spot rate between the U.S. dollar and the New Zealand dollar is $1 = NZD1.3352. Assume the interest rate in the United States is 4 percent and in New Zealand is 2 percent.

What should be the 6-month forward exchange rate?
Answer #1
6-month forward exchange rate = Spot rate*(1+p)/(1+b)
= 1.3352*(1+0.01)/(1+0.02)
=      1.3221
Where,
p 6 months interest rate of New Zealand = 2%*6/12 = 0.01
b 6 months interest rate of United States = 4%*6/12 = 0.02
Thus,
6 months forward exchange rate is
$             1 = NZD1.3221